OpenStudy (anonymous):

A pole is fixed vertically on a plane. An ant while walking on the plane, observes that its angle of elevation to the top of the pole is 45 degrees. Then it walks 1 m straight towards the pole, turns left and then moves 2 m further. If it finds the angle of elevation is again 45 degrees, then the height of the pole is a) 2.5 m b) 2 m c) root 2 m d) (root 3)/2 m

OpenStudy (anonymous):

 Therefore x^2 = (x-1)^2 + 2^2 4 = x + x - 1 Therefore x = 5/2